# Three Fundamental Questions in Modern Infinite-Domain Constraint Satisfaction

### v2

@@ 5.2. Our sandwich recipe
    ${\mathfrak{A}}^{[d]}$ is a cheese for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$;
@@
    Every finite cheese for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ pp-constructs $\mathfrak{K}_{3}$.
@@
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have $\smash{{\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$. Now suppose that there is a finite cheese $\mathfrak{D}$ of $({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\coloneqq h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ of arity $n$. One can show that every polymorphism from ${\mathfrak{S}}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ can be lifted to a polymorphism from $\mathfrak{S}$ to $\mathfrak{A}$ and that any height-1 identity that is satisfied in $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ is also satisfied in $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ modulo $\mathrm{Aut}(\mathfrak{B})$. Hence, there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. One easily sees that any operation that is pseudo-cyclic modulo some order-preserving automorphism group must be cyclic. This means that we have successfully lifted the cyclic identity from $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ to $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$. But any cyclic operation on a set of size at least 3 does not preserve $\neq$, a contradiction. ∎
@@
[…]
@@ 6.1. Topology seems irrelevant
In the context of infinite-domain CSPs, the influence of topological properties of the polymorphism clones of $\omega$-categorical structures has received much attention, with various results claiming their relevance [27] or irrelevance [11]: while generally, whether or not an $\omega$-categorical structure $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ (or pp-interprets another $\omega$-categorical structure $\mathfrak{A^{\prime}}$) does depend on such topological properties [9, 30, 46, 32], in many situations it does not (see, e.g., [41]). By moving from any $\omega$-categorical model-complete core template to one without algebraicity, Theorem 2.2 allows us to restrict Conjecture 1.4 to a class of topologically well-behaved structures. Namely, for those recent research suggests that often the algebraic structure determines the topological one; at least this is a fact for the endomorphism monoid, where the relevant topology (pointwise convergence) can then be defined from the algebraic structure of the monoid alone [62]. For $\omega$-categorical structures with algebraicity, on the other hand, various examples [42, 18, 62] show that this need not be the case. This also implies that there is no hope of lifting isomorphisms between endomorphism monoids or polymorphism clones of $\omega$-categorical structures to their counterparts constructed in Theorem 2.2.
@@ 6.2. Necessity of cheese preprocessing
We claim that $\mathrm{CSP}({\mathfrak{A}}^{[d]}_{S})$ is tractable, and hence this PCSP is finitely tractable. To see this, observe that the binary minimum operation $\min(x,y)$ is a polymorphism of $\mathfrak{A}$. This operation is conservative: $\min(x,y)\subseteq\{x,y\}$ holds for all $x,y$. Hence, its restriction to $A_{S}^{2}$ induces a polymorphism of $\mathfrak{A}_{S}$, which itself induces a polymorphism of $\smash{{\mathfrak{A}}^{[d]}_{S}}$ through its component-wise action. Since this operation is cyclic, $\smash{{\mathfrak{A}}^{[d]}_{S}}$ does not pp-construct $\mathfrak{K}_{3}$ [9, Theorem 1.4], and hence we are done by Theorem 1.1. This issue cannot be fixed simply by taking an expansion of $\mathfrak{A}$ by the inequality relation $\neq$ because $\mathrm{CSP}(\mathfrak{A};\neq)$ is NP-complete [22].
@@ 6.3. Algorithmic properties of our sandwiches
[…]
@@ A.1. Omitted definitions
$(x\approx y)\Leftarrow$ $I_{4}(u,v,x,y)\wedge(u\approx v),$
$(x\approx z)\Leftarrow$ $(x\approx y)\wedge(y\approx z),$
$(x\approx y)\Leftarrow$ $(y\approx x),$
$(x\approx y)\Leftarrow$ $(x=y).$
[…]
Suppose that $\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Note that $(\mathfrak{J}/{\approx})|^{\tau}$ is a substructure of $\mathcal{I}^{\ast}(\mathfrak{J})$, and hence $(\mathfrak{J}/{\approx})|^{\tau}\rightarrow\mathfrak{A}$. Since $\mathfrak{A}$ is CSP-injective, there exists an injective homomorphism $h\colon(\mathfrak{J}/{\approx})|^{\tau}\rightarrow\mathfrak{A}$. By the definition of $\approx$, if $([s_{1}]_{\approx},[s_{2}]_{\approx},[s_{3}]_{\approx},[s_{4}]_{\approx})\in I_{4}^{\mathfrak{J}/{\approx}}$ and $[s_{1}]_{\approx}=[s_{2}]_{\approx}$, then $[s_{3}]_{\approx}=[s_{4}]_{\approx}$. Since $h$ is injective, it follows that $h$ is also a homomorphism from $\mathfrak{J}/{\approx}$ to $\mathfrak{A}_{I_{4}}$. Then $h\circ q_{\approx}$ is a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}_{I_{4}}$.
@@
[…]
@@
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. By Example 1.3 and Example 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$ from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the proofs of CSP/Pol-injectivity and Datalog-interreducibility, starting with the CSP-injective case.
@@
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in ×$\mathbb{Q}\times\{a\}$ since this does not affect the relations in $\tau\cup\{E\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
@@
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$: otherwise, the map on $\mathfrak{A}$ sending every element to $h(q_{\sim}(x))$ would provide such constant endomorphism. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
Suppose that $e^{\ast}$ is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Let $S^{\ast}$ be an arbitrary finite subset of ×$\mathbb{Q}\times A$. Since $e^{\ast}$ preserves $\neq$, it is injective. Since $e^{\ast}$ preserves $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$, which is an equivalence relation, it naturally acts on its classes. Let $e$ be the map on (×$(\mathbb{Q}\times A)/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ mapping each $[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $[e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$, and let $S\coloneqq\{[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}\mid x\in S^{\ast}\}$. Similarly as in Proposition 4.5(1), we have that $e$ is an endomorphism of $\mathfrak{A}$. Since $\mathfrak{A}$ is a model-complete core, there exists an automorphism $\alpha$ of $\mathfrak{A}$ such that $\alpha|_{S}=e|_{S}$. We claim that there exists an automorphism ${\alpha}^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. To see this, note that we can simply fix a permutation of each equivalence class of $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ and extend $\alpha$, which is a permutation of the equivalence classes, according to it. Since $S^{\ast}$ is finite, we can moreover select the permutations so that ${\alpha}^{\ast}$ coincides with $e^{\ast}$ on $S^{\ast}$.
Next, suppose that $e$ is an endomorphism of $\mathfrak{A}$, and let $S$ be an arbitrary finite subset of $A$. Define $S^{\ast}\coloneqq\{0\}\times S$. Let $e^{\ast}$ be the map on ×$\mathbb{Q}\times A$ defined via $e^{\ast}(q,a)\coloneqq(q,e(a))$. This function is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is a model-complete core, there exists an automorphism ${\alpha}^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. Let $\alpha$ be the map on $A$ mapping each $[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $[e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ As before, we have that $\alpha$ is a well-defined automorphism of $\mathfrak{A}$ extending $e$ on $S$.

### v3

@@ 5.2. Our sandwich recipe
[…]
@@
[…]
@@
[…]
@@
Let $\mathfrak{A}_{1},\mathfrak{A}_{2}$ be countable homogeneous structures without algebraicity over disjoint finite relational signatures $\tau_{1},\tau_{2}$. Then there exists an up to isomorphism unique countable homogeneous $(\tau_{1}\cup\tau_{2})$-structure $\mathfrak{A}_{1}\smash{\raisebox{-6.64433pt}{{\LARGE\hskip 0.79253pt*\hskip 0.79253pt}}}\mathfrak{A}_{2}$, called the generic superposition of $\mathfrak{A}_{1},\mathfrak{A}_{2}$, such that any finite $(\tau_{1}\cup\tau_{2})$-structure embeds into $\mathfrak{A}_{1}\smash{\raisebox{-6.64433pt}{{\LARGE\hskip 0.79253pt*\hskip 0.79253pt}}}\mathfrak{A}_{2}$ if and only if its $\tau_{i}$-reduct embeds into $\mathfrak{A}_{i}$ for all $i\in\{1,2\}$. Moreover, if $\mathfrak{A}_{1},\mathfrak{A}_{2}$ are both finitely bounded, then so is $\mathfrak{A}_{1}\smash{\raisebox{-6.64433pt}{{\LARGE\hskip 0.79253pt*\hskip 0.79253pt}}}\mathfrak{A}_{2}$.
@@ 6.1. Topology seems irrelevant
In the context of infinite-domain CSPs, the influence of topological properties of the polymorphism clones of $\omega$-categorical structures has received much attention, with various results claiming their relevance [27] or irrelevance [11]: while generally, whether or not an $\omega$-categorical structure $\mathfrak{A}$ $p$-constructs $\mathfrak{K}_{3}$ (or pp-interprets another $\omega$-categorical structure $\mathfrak{A^{\prime}}$) does depend on such topological properties [9, 30, 46, 32], in many situations it does not (see, e.g., [41]). By moving from any $\omega$-categorical model-complete core template to one without algebraicity, Theorem 2.2 allows us to restrict Conjecture 1.4 to a class of topologically well-behaved structures. Namely, for those recent research suggests that often the algebraic structure determines the topological one; at least this is a fact for the endomorphism monoid, where the relevant topology (pointwise convergence) can then be defined from the algebraic structure of the monoid alone [62]. For $\omega$-categorical structures with algebraicity, on the other hand, various examples [42, 18, 62] show that this need not be the case. This also implies that there is no hope of lifting isomorphisms between endomorphism monoids or polymorphism clones of $\omega$-categorical structures to their counterparts constructed in Theorem 2.2.
@@ 6.2. Necessity of cheese preprocessing
[…]
@@ 6.3. Algorithmic properties of our sandwiches
Open question: Is there a universal algorithm that would solve all tractable PCSPs stemming from Theorem 2.6, even when $\widehat{\mathfrak{A}}=(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is replaced by $\widehat{\mathfrak{A}}=\big{(}\mathbb{Q}\wr\mathfrak{A}\big{)}_{\neq,I_{4}}$?
@@ A.1. Omitted definitions
$(x\approx y)\Leftarrow\$$I_{4}(u,v,x,y)\wedge(u\approx v),$
$(x\approx z)\Leftarrow\$$(x\approx y)\wedge(y\approx z),$
$(x\approx y)\Leftarrow\$$(y\approx x),$
$(x\approx y)\Leftarrow\$$(x=y).$
By definition, this relation is an equivalence relation on $J$; denote by $q_{\approx}$ be the induced quotient map $J\rightarrow J/{\approx},x\mapsto[x]_{\approx}$. Let $S\coloneqq\{s_{1},\dots,s_{n}\}$ be arbitrary representatives of the $\approx$}-classes. Without loss of generality, $S=J/{\approx}$. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\approx}$ under $q_{\approx}$. We verify that $\mathcal{I}^{\ast}$ is a reduction.
Suppose that $\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Note that $(\mathfrak{J}/{\approx})|^{\tau}$ is a substructure of $\mathcal{I}^{\ast}(\mathfrak{J})$, and hence $(\mathfrak{J}/{\approx})|^{\tau}\rightarrow\mathfrak{A}$. Since $\mathfrak{A}$ is CSP-injective, there exists an injective homomorphism $h\colon(\mathfrak{J}/{\approx})|^{\tau}\rightarrow\mathfrak{A}$. By the definition of $\approx$, if $([s_{1}]_{\approx},[s_{2}]_{\approx},[s_{3}]_{\approx},[s_{4}]_{\approx})\in I_{4}^{\mathfrak{J}/{\approx}}$ and $s_{1}]_{\approx}=[s_{2}]_{\approx}$, then $s_{3}]_{\approx}=[s_{4}]_{\approx}$. Since $h$ is injective, it follows that $h$ is also a homomorphism from $\mathfrak{J}/{\approx}$ to $\mathfrak{A}_{I_{4}}$. Then $h\circ q_{\approx}$ is a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}_{I_{4}}$.
@@
which is a union of basic open sets. By Lemma 6.7(iii) in [52], products of extremely amenable groups are extremely amenable, and hence $\mathrm{Aut}(\mathfrak{A})^{B}$ is extremely amenable. If we can show that also $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})/\mathrm{Aut}(\mathfrak{A})^{B}$ is extremely amenable, then it follows from Lemma 6.7(ii) in [52] that $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$ is extremely amenable as well. The composition $\pi\circ\iota$ of the natural embedding $\iota\colon\mathrm{Aut}(\mathfrak{B})\rightarrow\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$ with the natural quotient map $\pi\colon\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})\rightarrow\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})/\mathrm{Aut}(\mathfrak{A})^{B}$ is a topological isomorphism between $\mathrm{Aut}(\mathfrak{B})$ and $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})/\mathrm{Aut}(\mathfrak{A})^{B}$. Hence, the claim follows from the fact that $\mathrm{Aut}(\mathfrak{B})$ is extremely amenable. Recall that, by item 4, we have that $\mathfrak{A}\wr\mathfrak{B}$ is homogeneous. Since $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$ is extremely amenable, we conclude using Theorem A.1 that $\mathfrak{A}\wr\mathfrak{B}$ is homogeneous Ramsey. ∎
@@
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. By Example 1.3 and Example 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol << from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the proofs of CSP/Pol-injectivity and Datalog-interreducibility, starting with the CSP-injective case.
@@
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in $\mathbb{Q}\times\{a\}$ since this does not affect the relations in $\tau\cup\{E\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
@@
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$: otherwise, the map on $\mathfrak{A}$ sending every element to $h(q_{\sim}(x))$ would provide such constant endomorphism. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
Suppose that $e^{\ast}$ is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Let $S^{\ast}$ be an arbitrary finite subset of $\mathbb{Q}\times A$. Since $e^{\ast}$ preserves $\neq$, it is injective. Since $e^{\ast}$ preserves $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$, which is an equivalence relation, it naturally acts on its classes. Let $e$ be the map on $(\mathbb{Q}\times A)/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ mapping each $x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$, and let $S\coloneqq\{[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}\mid x\in S^{\ast}\}$. Similarly as in Proposition 4.5(1), we have that $e$ is an endomorphism of $\mathfrak{A}$. Since $\mathfrak{A}$ is a model-complete core, there exists an automorphism $\alpha$ of $\mathfrak{A}$ such that $\alpha|_{S}=e|_{S}$. We claim that there exists an automorphism $\alpha$$^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. To see this, note that we can simply fix a permutation of each equivalence class of $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ and extend $\alpha$, which is a permutation of the equivalence classes, according to it. Since $S^{\ast}$ is finite, we can moreover select the permutations so that $\alpha$$^{\ast}$ coincides with $e^{\ast}$ on $S^{\ast}$.
Next, suppose that $e$ is an endomorphism of $\mathfrak{A}$, and let $S$ be an arbitrary finite subset of $A$. Define $S^{\ast}\coloneqq\{0\}\times S$. Let $e^{\ast}$ be the map on $\mathbb{Q}\times A$ defined via $e^{\ast}(q,a)\coloneqq(q,e(a))$. This function is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is a model-complete core, there exists an automorphism $\alpha$$^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. Let $\alpha$ be the map on $A$ mapping each $x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ As before, we have that $\alpha$ is a well-defined automorphism of $\mathfrak{A}$ extending $e$ on $S$.